# Diagnostically Competitive Performance of a Physiology-Informed Generative Multi-Task Network for Contrast-Free CT Perfusion

### v1

@@ 4 Discussion
The proposed MAGIC model, including the novel physiology-informed loss terms and physicians-in-the-loop module, performs favorably in comparison to commercial software for perfusion imaging with contrast. MAGIC enables rapid, contrast-free, high fidelity, high resolution, and physician-preferred generation of CTP imaging from only the NCCT series. Implementing this model in clinical practice has the potential to: reduce the need for contrast injection, produce diagnostically equivalent perfusion imaging, reducing acquisition time (roughly 90 seconds or more using contrast-enhanced CTP [59] to roughly 1.5 seconds using NCCT only [60]), forgoing the length of preparation time for contrast injection, improving scan-to-map post-processing speed (roughly 5 minutes using RAPID software 31 to about 7.43 sec per volume of 161 slices once trained), providing cost-effectiveness at which physicians can assess a patient’s hemodynamic activity, and saving critical time at rapid triage stage (Fig. 7)[40]. This can further be used to enable CTP imaging in limited-resource hospitals that currently lack the proper training, staff, contrast agent, funds, or facilities to conduct CTP protocols. Additionally, the use of non-contrast imaging in synthesizing perfusion data via MAGIC presents potential benefits in reducing the radiation dose of routine imaging procedures. The radiation output of a head CTP exam is roughly 4 to 10 times that of a head NCCT exam (200 to 500 mGy in CTP vs. 50 to 70 mGy in NCCT). Prior studies have raised the concern of patient harm from radiation exposure during CTP, especially within the context of multimodal CT protocol[61, 62]. The use of non-contrast imaging in the synthesis of CTP imaging also circumvents the potential negative side effects associated with iodinated contrast injection including: allergic reactions, cardiac and renal failure, and pulmonary edema [63]. Additionally, the elimination of intravenous contrast injection can significantly reduce the supply cost. These improvements to the safety, efficiency, cost, and accessibility of CTP imaging positions MAGIC as a valuable asset in enabling contrast-free perfusion imaging.
@@
Figure 7: Proposed advantages of MAGIC and comparison against existing workflows. a. The amount of time from arriving to acquiring CT perfusion maps is significantly reduced. Preparation and scan acquisition for CT perfusion is cut out from the MAGIC pipeline. Times are estimated from existing workflows. b. MAGIC removes the preparation and resources necessary to acquire CT perfusion maps. This pipeline reduces the number of CT perfusion contraindications. Removal of contrast injection allows for rapid triage for stroke treatment, faster CT perfusion acquisition with only NCCT, reduced monetary cost, minimized radiation exposure, and reduced preparation time.
@@
This paper is the first to address the research question and clinical potential of a contrast-free generation of CTP imaging, using image-to-image translation techniques between the NCCT and CTP series. We introduced significant technical and clinical innovations into this physiology-informed multitask deep learning model. The novel physicians-in-the-loop (PILO) module provides radiologists with an interactive and adaptive diagnostic tool to select and visualize the hemodynamic activity embedded in anatomical imaging. The radiologist evaluators in this work stated that: PILO is a valuable tool that allowed for better visualization and diagnosis of the brain’s perfusion activity than the clinical perfusion scans alone.
@@
Our evaluation demonstrates that MAGIC can utilize deep learning-based multimodal image translation techniques to synthesize non-contrast perfusion maps. Our synthetic perfusion imaging has comparable structural fidelity and diagnostic quality to real perfusion imaging generated using commercial software. Perfusion maps generated by MAGIC, especially CBF and CBV maps, displayed accurate hemodynamic status present in real imaging. There was no significant difference in the values of the image metrics across all four image map classes, which indicates that MAGIC performs stably across perfusion map types. Additionally, there was no significant difference between the performance metrics across different patients with varying levels of infarction status, which indicates that MAGIC is robust to abnormal hemodynamic activities such as acute ischemia.
@@
Our double-blinded study revealed that our synthetic perfusion imaging accurately portrays real imaging. Additionally, our synthetic perfusion imaging has no distinguishable features that the clinicians could use to consistently discriminate between synthesized and real imaging. The clinicians evaluators in this study concluded that the synthetic imaging was of sufficient diagnostic quality to confidently diagnose a patient and generate a treatment plan. In concurrent discussions with the clinicians, we found that they prefer to use CBV and TTP imaging to produce a diagnostic decision and characterize the volumes of the infarcted core and the ischemic penumbra. We found that the diagnoses and treatment plans the clinicians produced using the synthesized imaging were statistically consistent with those they gave to the corresponding real perfusion imaging.
@@
The diagnostic accuracy and computational efficiency offered by MAGIC potentiates more rapid and cost-efficient clinical workflows in assessing neurovascular health and treating acute ischemia. Meanwhile, there are some limitations of this study. Since non-contrast CTP synthesis via deep learning image translation is a novel research problem, there are no baseline approaches to which we can compare our results. Additionally, we found that the MTT imaging produced by the RAPID software was consistently ranked as “unacceptable” when evaluated by clinicians. The RAPID data have limitations that account for its current limited use in the clinical setting. MAGIC is unable to capture a complete representation of microvascular activity consistently, even though it performs well in predicting hemodynamic activity from non-contrast imaging. For example, the current MAGIC iteration cannot predict small-vessel occlusions based on non-contrast imaging alone (Supplementary Fig. 7). This limitation is evident in cases where low-density areas in the periventricular white matter, typically indicative of small-vessel ischemic injury, are not captured by the perfusion parameters. We plan to address these issues in future improvements of MAGIC by incorporating more detailed perfusion imaging data from other cohorts into the training set. Furthermore, we aim to address the issues present with capturing the activity from the temporal perfusion maps, TTP and MTT, through further iterations on our methodology and diversity of our training set. A more accurate encoding of the temporal map classes will enable clinicians to make a more informed diagnosis. While the MAGIC model has shown promising results, its generalizability to external datasets remains a key area for future work. It has not yet been tested on external data, which may vary in clinical settings, imaging protocols, and patient demographics. Evaluating the model on diverse, independent datasets is essential for assessing its robustness and clinical applicability. Additionally, to improve synthetic CTP map generation, we plan to incorporate advanced models like diffusion models, which could enhance the accuracy and consistency of temporal maps and improve microvascular activity predictions, ultimately increasing the clinical utility of the generated maps [64, 65, 66, 67, 68]. Finally, NCCT and CTP slices in the training or test sets may not be registered perfectly due to patient movement between scanning protocols while collecting imaging data. Our training framework requires paired and co-registered data between the NCCT and CTP imaging series. Considering this, we must only use the NCCT image slices with a corresponding CTP slice at the exact vertical location, since the axial CTP image slices are sparser in spatial density than the axial NCCT image slices. As a result, we could not incorporate our entire dataset into our training algorithm to fully leverage information from all slices in NCCT. MAGIC is still capable of learning the mapping between NCCT and CTP well and its ability to generalize to new data, even with the reduced training data size.
@@ Acknowledgements
Supplementary Materials
Supplementary Material
Figure S1: A diagram of the generator network architecture. The generator follows a modified U-Net architecture, with various network layers connected by skip connections. The first encoding layers share the encoded weights of each map before diverging into multi-task perfusion map generation. The physicians-in-the-loop module comprises the final layers of the architecture in which the NCCT input is concatenated with the initial perfusion output. A final deconvolutional layer is applied to generate the final perfusion map outputs.
Figure S2: Discriminator network architecture which is based on the PatchGAN framework, which evaluates the realism of local image patches rather than entire images. The discriminator applies a series of transposed convolutional layers to the concatenated input of the NCCT slice and the CTP slice for evaluation.
Figure S3: The questionnaire provided to evaluators for double-blinded evaluation.
Table S1: The results of Q1 from the double-blinded questionnaire: “Do you think this is a real CT perfusion map?”. All raters were unable to differentiate between real and synthetic CTP imaging. Overall, both real (RAPID) and synthetic (MAGIC) CTP images were predominantly classified as "RAPID"— 117 and 118 responses, respectively. This trend held across individual doctors, with the exception of Doctor 7, who more frequently identified images as MAGIC. These findings suggest that raters predominantly treated both real and synthetic CTP maps as real, indicating that MAGIC is capable of generating synthetic CTP maps that closely resemble real ones.
Doctor RAPID MAGIC
Is RAPID Is MAGIC Is RAPID Is MAGIC
1 12 8 15 5
2 19 1 19 1
3 19 1 16 4
4 20 0 18 2
5 20 0 18 2
6 19 1 19 1
7 8 12 13 7
Sum 117 23 118 22
Table S2: Comparison of individual responses to RAPID and MAGIC CTP for Questions 2: “What is the diagnostic quality of this perfusion map?” The table shows the response counts (Acceptable, Indeterminate, Unacceptable) for CBF maps processed by RAPID and MAGIC. Most images were rated “Acceptable” across both methods, with very few “Indeterminate” or “Unacceptable” responses. Acc = Acceptable, Indet = Indeterminate, Unacc = Unacceptable.
Doctor RAPID CBF MAGIC CBF
Acc Indet Unacc Acc Indet Unacc
1 20 0 0 19 1 0
2 20 0 0 19 1 0
3 20 0 0 20 0 0
4 20 0 0 18 2 0
5 20 0 0 17 3 0
6 20 0 0 20 0 0
7 20 0 0 20 0 0
Sum 140 0 0 133 7 0
Table S3: Response comparison for RAPID and MAGIC CTP on Question 2: “What is the diagnostic quality of this perfusion map?” Counts of Acceptable (Acc), Indeterminate (Indet), and Unacceptable (Unacc) ratings for CBV maps are shown. Most images were rated “Acceptable” for both methods.
Doctor RAPID CBV MAGIC CBV
Acc Indet Unacc Acc Indet Unacc
1 20 0 0 19 1 0
2 20 0 0 18 1 1
3 18 2 0 18 1 1
4 20 0 0 18 2 0
5 20 0 0 17 1 2
6 20 0 0 20 0 0
7 20 0 0 19 0 1
Sum 138 2 0 129 6 5
Table S4: Comparison of responses for RAPID and MAGIC CTP on Question 2: Counts of Acceptable (Acc), Indeterminate (Indet), and Unacceptable (Unacc) ratings for MTT maps. Most images were rated “Acceptable” for both methods.
Doctor RAPID MTT MAGIC MTT
Acc Indet Unacc Acc Indet Unacc
1 20 0 0 20 0 0
2 14 0 6 7 4 9
3 18 2 0 19 1 0
4 4 2 14 0 2 18
5 20 0 0 18 2 0
6 20 0 0 20 0 0
7 17 0 3 17 0 3
Sum 113 4 23 101 9 30
Table S5: Comparison of responses for RAPID and MAGIC CTP on Question 2: Counts of Acceptable (Acc), Indeterminate (Indet), and Unacceptable (Unacc) ratings for TTP maps. Most images were rated “Acceptable” with few “Indeterminate” or “Unacceptable” responses.
Doctor RAPID TTP MAGIC TTP
Acc Indet Unacc Acc Indet Unacc
1 19 1 0 19 1 0
2 18 2 0 15 4 1
3 20 0 0 18 2 0
4 14 5 1 14 6 0
5 20 0 0 18 2 0
6 20 0 0 20 0 0
7 18 0 2 20 0 0
Sum 129 8 3 124 15 1
Table S6: Results for responses to Q3: “What diagnosis would be given to this patient based on their NCCT and CTP imaging?” The matrix on the left shows the distribution of diagnostic ratings (1–4) assigned to each case using real CTP (RAPID) versus synthetic CTP (MAGIC). The majority of responses are concentrated along the diagonal, indicating high agreement between diagnoses made with real and synthetic CTP images. For example, 43 cases were rated as category 1 by both methods, and similar agreement is seen for other categories.
Method MAGIC
Option 1 2 3 4
RAPID 1 43 14 0 2
2 13 9 2 6
3 3 2 6 13
4 6 2 1 16
Table S7: Results for responses to Q4: “Rate your confidence in the above diagnosis.” The matrix on the right displays the confidence ratings (1–5) for each case, comparing real and synthetic CTP. Most responses are clustered along the diagonal, especially at higher confidence levels (e.g., 44 cases rated as 4 by both methods, and 25 cases rated as 5 by both), indicating that raters’ confidence in their diagnoses was generally consistent between RAPID and MAGIC CTP images.
Method MAGIC
Option 1 2 3 4 5
RAPID 1 1 0 1 0 0
2 0 3 1 4 2
3 0 1 4 8 1
4 0 9 6 44 15
5 0 1 3 11 25
Table S8: Ablation study results for the MAGIC model across four perfusion maps: Cerebral Blood Flow (CBF), Cerebral Blood Volume (CBV), Mean Transit Time (MTT), and Time to Peak (TTP). SSIM and UQI scores are reported for the full MAGIC model and for variants with specific loss components removed (no multimodal loss, no extrema loss, and no loss components). The results highlight the importance of each loss in contributing to the final performance.
Model CBF CBV MTT TTP
SSIM UQI SSIM UQI SSIM UQI SSIM UQI
No-multimodal 0.7911 0.6129 0.7861 0.6222 0.8103 0.7422 0.79716 0.7203
No-both 0.7936 0.6186 0.7886 0.6243 0.8130 0.7442 0.80007 0.7234
No extrema 0.7835 0.6148 0.7807 0.5454 0.8036 0.7529 0.7997 0.7739
MAGIC 0.8302 0.9313 0.8158 0.9037 0.7938 0.9453 0.7975 0.9066
Table S9: Comparison of MAGIC with state-of-the-art models (U-Net and Pix2Pix) across CBF, CBV, MTT, and TTP.
Model CBF CBV MTT TTP
SSIM UQI SSIM UQI SSIM UQI SSIM UQI
UNet 0.6164 0.7747 0.7951 0.9093 0.8308 0.8783 0.8105 0.8600
Pix2pix 0.7561 0.4519 0.8068 0.8828 0.8320 0.8353 0.8176 0.8152
MAGIC 0.8302 0.9313 0.8158 0.9037 0.7938 0.9453 0.7975 0.9066
Figure S4: Example of small-vessel ischemic injury on non-contrast CT imaging. This figure demonstrates the substantial low density of the periventricular white matter, especially near the occipital horns, which is typically indicative of small vessel ischemic injury. Despite this, the perfusion parameters measured in the current MAGIC iteration do not reflect this condition, highlighting a limitation of the proposed method in detecting small-vessel occlusions based solely on non-contrast imaging.

### v2

@@ 4 Discussion
The proposed MAGIC model, including the novel physiology-informed loss terms and physicians-in-the-loop module, performs favorably in comparison to commercial software for perfusion imaging with contrast. MAGIC enables rapid, contrast-free, high fidelity, high resolution, and physician-preferred generation of CTP imaging from only the NCCT series. Implementing this model in clinical practice has the potential to: reduce the need for contrast injection, produce diagnostically comparable perfusion imaging, reducing acquisition time (roughly 90 seconds or more using contrast-enhanced CTP [20] to roughly 1.5 seconds using NCCT only [54]), forgoing the length of preparation time for contrast injection, improving scan-to-map post-processing speed (roughly 5 minutes using RAPID software 31 to about 7.43 sec per volume of 161 slices once trained), providing cost-effectiveness at which physicians can assess a patient’s hemodynamic activity, and saving critical time at rapid triage stage (Fig. 7)[44]. This can further be used to enable CTP imaging in limited-resource hospitals that currently lack the proper training, staff, contrast agent, funds, or facilities to conduct CTP protocols. Additionally, the use of non-contrast imaging in synthesizing perfusion data via MAGIC presents potential benefits in reducing the radiation dose of routine imaging procedures. The radiation output of a head CTP exam is roughly 4 to 10 times that of a head NCCT exam (200 to 500 mGy in CTP vs. 50 to 70 mGy in NCCT). Prior studies have raised the concern of patient harm from radiation exposure during CTP, especially within the context of multimodal CT protocol[67, 31]. The use of non-contrast imaging in the synthesis of CTP imaging also circumvents the potential negative side effects associated with iodinated contrast injection including: allergic reactions, cardiac and renal failure, and pulmonary edema [6]. Additionally, the elimination of intravenous contrast injection can substantially reduce the supply cost. These improvements to the safety, efficiency, cost, and accessibility of CTP imaging positions MAGIC as a valuable asset in enabling contrast-free perfusion imaging.
@@
Figure 7: Proposed advantages of MAGIC and comparison against existing workflows. a. The amount of time from arriving to acquiring CT perfusion maps is considerably reduced. Preparation and scan acquisition for CT perfusion is cut out from the MAGIC pipeline. Times are estimated from existing workflows. b. MAGIC removes the preparation and resources necessary to acquire CT perfusion maps. This pipeline reduces the number of CT perfusion contraindications. Removal of contrast injection allows for rapid triage for stroke treatment, faster CT perfusion acquisition with only NCCT, reduced monetary cost, minimized radiation exposure, and reduced preparation time.
@@
This paper is the first to address the research question and clinical potential of a contrast-free generation of CTP imaging, using image-to-image translation techniques between the NCCT and CTP series. We introduced cutting edge innovations into this physiology-informed multitask deep learning model. The novel physicians-in-the-loop (PILO) module provides radiologists with an interactive and adaptive diagnostic tool to select and visualize the hemodynamic activity embedded in anatomical imaging. The radiologist evaluators in this work stated that: PILO is a valuable tool that allowed for better visualization and diagnosis of the brain’s perfusion activity than the clinical perfusion scans alone.
@@
Our evaluation demonstrates that MAGIC can utilize deep learning-based multimodal image translation techniques to synthesize non-contrast perfusion maps. Our synthetic perfusion imaging has comparable structural fidelity and diagnostic quality to real perfusion imaging generated using commercial software. Perfusion maps generated by MAGIC, especially CBF and CBV maps, displayed accurate hemodynamic status present in real imaging. There was no meaningful difference in the values of the image metrics across all four image map classes, which indicates that MAGIC performs stably across perfusion map types. Additionally, there was no meaningful difference between the performance metrics across different patients with varying levels of infarction status, which indicates that MAGIC is robust to abnormal hemodynamic activities such as acute ischemia.
@@
Our double-blinded study revealed that our synthetic perfusion imaging accurately portrays real imaging. Additionally, our synthetic perfusion imaging has no distinguishable features that the clinicians could use to consistently discriminate between synthesized and real imaging. The clinicians evaluators in this study concluded that the synthetic imaging was of sufficient diagnostic quality to confidently diagnose a patient and generate a treatment plan. In concurrent discussions with the clinicians, we found that they prefer to use CBV and TTP imaging to produce a diagnostic decision and characterize the volumes of the infarcted core and the ischemic penumbra. We found that the diagnoses and treatment plans the clinicians produced using the synthesized imaging were statistically consistent with those they gave to the corresponding real perfusion imaging. Furthermore, although a mixed-effects analysis could further quantify reader-level variability, the current sample of ratings per doctor was not sufficient to support a stable generalized linear mixed model; therefore, we relied on descriptive agreement patterns for this study.
@@
The diagnostic accuracy and computational efficiency offered by MAGIC potentiates more rapid and cost-efficient clinical workflows in assessing neurovascular health and treating acute ischemia. Meanwhile, there are some limitations of this study. Since non-contrast CTP synthesis via deep learning image translation is a novel research problem, there are no baseline approaches to which we can compare our results. Additionally, we found that the MTT imaging produced by the RAPID software was consistently ranked as “unacceptable” when evaluated by clinicians. The RAPID data have limitations that account for its current limited use in the clinical setting. MAGIC is unable to capture a complete representation of microvascular activity consistently, even though it performs well in predicting hemodynamic activity from non-contrast imaging. For example, the current MAGIC iteration cannot predict small-vessel occlusions based on non-contrast imaging alone (Supplementary Fig. 7). This limitation is evident in cases where low-density areas in the periventricular white matter, typically indicative of small-vessel ischemic injury, are not captured by the perfusion parameters. We plan to address these issues in future improvements of MAGIC by incorporating more detailed perfusion imaging data from other cohorts into the training set.
Furthermore, we aim to address the issues present with capturing the activity from the temporal perfusion maps, TTP and MTT, through further iterations on our methodology and diversity of our training set. A more accurate encoding of the temporal map classes will enable clinicians to make a more informed diagnosis. While the MAGIC model has shown promising results, its generalizability to external datasets remains a key area for future work. It has not yet been tested on external data, which may vary in clinical settings, imaging protocols, and patient demographics. Evaluating the model on diverse, independent datasets is essential for assessing its robustness and clinical applicability. Additionally, to improve synthetic CTP map generation, we plan to incorporate advanced models like diffusion models, which could enhance the accuracy and consistency of temporal maps and improve microvascular activity predictions, ultimately increasing the clinical utility of the generated maps [47, 2, 22, 23, 58]. Furthermore, NCCT and CTP slices in the training or test sets may not be registered perfectly due to patient movement between scanning protocols while collecting imaging data. Our training framework requires paired and co-registered data between the NCCT and CTP imaging series. Considering this, we must only use the NCCT image slices with a corresponding CTP slice at the exact vertical location, since the axial CTP image slices are sparser in spatial density than the axial NCCT image slices. As a result, we could not incorporate our entire dataset into our training algorithm to fully leverage information from all slices in NCCT. MAGIC is still capable of learning the mapping between NCCT and CTP well and its ability to generalize to new data, even with the reduced training data size. Finally, although MAGIC can infer perfusion patterns from NCCT, it lacks interpretability to identify the NCCT features that drive these predictions. Therefore, MAGIC should be viewed as learning statistical associations rather than explicit physiological mechanisms.